Each of the following scenarios is independent. Assume that all cash flows are after-tax cash flows. a. Campbell Manufacturing is considering the purchase of a new welding system. The cash benefits will be $480,000 per year. The system costs $2,700,000 and will last 10 years. b. Evee Cardenas is interested in investing in a women's specialty shop. The cost of the investment is $270,000. She estimates that the return from owning her own shop will be $52,500 per year. She estimates that the shop will have a useful life of 6 years. c.Barker Company calculated the NPV of a project and found it to be $63,900. The project's life was estimated to be 8 years. The required rate of return used for the NPV calculation was 10%. The project was expected to produce annual after-tax cash flows of $135,000. Required: Compute the NPV or Campbell Manufacturing, assuming a discount rate of 12% f required round all present va 1. e calculation stone nearest dollar Should the company buy the new welding system? Yes 2. Conceptual Connection: Assuming a required rate of return of 8% calculate the N for Eve ardenas' investment. Round to the nearest dollar. re u red round all present value calculations to the nearest dollar. Use the minus sign to indicate a negative NPV Should she invest? No What if the estimated return was $135,000 per year? Calculate the new NPV for Evee Cardenas' investment. Would this affect the decision? What does this tell you about your analysis? Round to the nearest dollar. The shop should now be purchased. This reveals that the decision to accept or reject in this case is affected by differences in estimated cash flow
2. Conceptual Connection: Assuming a required rate of return of 8%, calculate the NPV for Evee Cardenas' investment. Round to the nearest dollar. If required, round all present value calculations to the nearest dollar. Use the minus sign to indicate a negative NPV Should she invest? No What if the estimated return was $135,000 per year? Calculate the new NPV for Evee Cardenas' investment. Would this affect the decision? What does this tell you about your analysis? Round to the nearest dollar The shop should now be purchased. This reveals that the decision to accept or reject in this case is affected by differences in estimated cash flow 3. What was the required investment for Barker Company's project? Round to the nearest dollar. If required, round all present value calculations to the nearest dollar
Answer #1
  1. Computation of NPV Campbel manufacturing

NPV = Present value of Cash Inflows - Present value of Cash Outflows.

PV OF CASH INFLOW = 480,000*5.650233= 2,712,107

PV OF CASH OUTFLOW=2,700,000

NPV= 2,712,107-2,700,000=12,107

Selection Criteria: If NPV > 0, i.e., if NPV is positive, the project is acceptable

YEAR

CASH INFLOW

PV @12%

PRESENT VALUE OF CASH INFLOW

1

480,000

1/1.12^1= 0.892857

428571

2

480,000

1/1.12^2 =0.797194

382653

3

480,000

0.71178

341655

4

480,000

0.635518

305049

5

480,000

0.567427

272365

6

480,000

0.506631

243183

7

480,000

0.452349

217128

8

480,000

0.403883

193864

9

480,000

0.36061

173093

10

480,000

0.321973

154547

TOTAL PRESENT VALUE OF CASH INFLOW

2,712,107

0

CASH OUTFLOW

(INITIAL INVESTMENT)

(2,700,000)

NET PRESENT VALUE

12,107

  1. Computation of NPV of Evee cardenas

NPV = Present value of Cash Inflows - Present value of Cash Outflows.

PV OF CASH INFLOW = 52,500* 4.62288= 242,701

PV OF CASH OUTFLOW=270,000

NPV= 242,701-270,000= (27299)

Selection Criteria: If NPV < 0, i.e., if NPV is negative, the project is not acceptable

If estimated annual cash inflow is $135,000

PV OF CASH INFLOW = 135,000* 4.62288= 624,089

PV OF CASH OUTFLOW=270,000

NPV= 624,089-270,000= 354,089

Selection Criteria: If NPV > 0, i.e., if NPV is positive, the project is acceptable